# ConsistencyTTA: Accelerating Diffusion-Based Text-to-Audio Generation with Consistency Distillation

### v2

@@ 1 Introduction
We focus on in-the-wild audio generation which produces a wide array of samples capturing the diversity of real-world sounds. Our extensive experiments, summarized in Figure 1, show that ConsistencyTTA simultaneously achieves high generation quality, fast inference speed, and high generation diversity. Specifically, the generation quality of the single-network-query ConsistencyTTA is comparable to a 400-query diffusion model across five objective metrics and two subjective metrics (audio quality and audio-text correspondence). Details explanations of Figure 1 are provided in Section A.1. We strongly encourage the reader to visit ConsistencyTTA’s demo at the website 111consistency-tta.github.io/demo.
@@
Using standard PyTorch implementation, ConsistencyTTA generates one minute of audio in just $9.1$ seconds on a laptop computer. In contrast, a representative diffusion method [1] requires over a minute on a state-of-the-art A100 GPU for an equivalent task (details in Section B.5).
@@ 2.1 Diffusion Models
Diffusion models [14, 15], known for their diverse and high-quality generations, have rapidly gained popularity across vision and audio generation tasks [10, 16, 3, 17, 18]. In vision, while pixel-level diffusion (e.g., EDM [16]) excel in generating small images, producing larger images requires LDMs [10] as they facilitate the diffusion process within a latent space. In the audio domain, while some works considered autoregressive models [8] or Mel-space diffusion [9], LDMs have emerged as the dominant TTA approach [1, 2, 3, 4, 5, 6, 7].
@@ 4.2 Main Evaluation Results
All diffusion baseline models use 200 inference steps following [3, 1], each step needing two noise estimations due to CFG, summing to 400 network queries per generation. Hence, we can conclude that with minimal performance drop, ConsistencyTTA reduces the U-Net queries by a factor of 400.
@@ 4.4 Audio Generation Diversity
ConsistencyTTA produces diverse generations as do diffusion models. Different random seeds (different initial Gaussian embeddings) produce noticeably different audio. To demonstrate, we present the generated waveforms from the first 50 AudioCaps test prompts with four different seeds at the website 222consistency-tta.github.io/diversity. We display the corresponding spectrograms, along with quantitative generation diversity analyses, in Section A.5.
@@ A.2 Real-World Inference Computing Time Comparison
On an Nvidia A100 GPU, generating from all 882 AudioCaps test prompts requires 2.3 minutes with our consistency model. The default TANGO model needs 168 minutes (73 minutes with the smaller 557M U-Net), 72 times as long compared with our consistency model. Note that the 200-step default inference schedule is shared among multiple diffusion-based TTA methods [1, 3], and thus, this TANGO inference time is representative. Moreover, our consistency model can run on a standard laptop computer, only taking 76 seconds to generate 50 ten-second audio clips, averaging 9.1 seconds per minute-generation. In contrast, the default TANGO requires 68 seconds per minute-generation on a state-of-the-art A100 GPU.
@@ A.5 More Generation Diversity Evidences
For quantitative evidence, we collect the Mel spectrograms of these 50 generations across four seeds, standardize them individually, calculate the standard deviation across different seeds, and average the deviations across all Mel spectrogram points of the 50 prompts. The average number is 0.871, again demonstrating non-trivial generation diversity.
@@ B.5 Evaluation Details
While the maximal audio length of the AudioCaps dataset is 10.00 seconds and the U-Net module of the TTA models is trained to generate 10.00-second latent audio representations, the HiFi-GAN vocoder produces 10.24-second audio. We observe that this mismatch negatively impacts the generation quality. Specifically, the final 0.24 seconds of the generated audio is empty, and there are slight distorting artifacts near the end of the 10-second useful portion. To this end, for the objective evaluation results in Tables 1 and 2, we truncate the generated audio to 9.70 seconds. Table 2 uses the full 10.24 seconds. The ground-truth reference waveforms are kept as-is. For $\text{CLAP}_{\text{A}}$ and $\text{CLAP}_{\text{T}}$ calculations, we use the CLAP checkpoint from [40] trained on LAION-Audio-630k [40], AudioSet [41], and music.
The human evaluation results in Table 1 are based on 20 evaluators each rating 25 audio clips per model, forming 500 samples per model. All AudioCaps captions are in English, and all evaluators are proficient in English, using it as their main business language. For each evaluator, the three models and the ground truth use the same set of prompts (the prompts vary across evaluators). Each evaluator rates each audio on a scale of 1 to 5, with rating criteria defined in the evaluation form. To ensure evaluation fairness, the model type generating each waveform is not disclosed to the evaluator, and the generations of the models are shuffled. We find it extremely challenging for a human to distinguish the outputs from the three generative models, with many ground truth waveforms also indistinguishable. An example evaluation form is available at consistency-tta.github.io/evaluation.

### v3

@@ 1 Introduction
We focus on in-the-wild audio generation which produces a wide array of samples capturing the diversity of real-world sounds. Our extensive experiments, summarized in Figure 1, show that ConsistencyTTA simultaneously achieves high generation quality, fast inference speed, and high generation diversity. Specifically, the generation quality of the single-network-query ConsistencyTTA is comparable to a 400-query diffusion model across five objective metrics and two subjective metrics (audio quality and audio-text correspondence). Detailed explanations of Figure 1 are provided in Section A.1.
@@
Using standard PyTorch implementation, ConsistencyTTA enables on-device audio generation, producing one minute of audio in only $9.1$ seconds on a laptop computer. In contrast, a representative diffusion method [1] requires over a minute on a state-of-the-art A100 GPU (see details in Section B.5).
We strongly encourage the reader to visit ConsistencyTTA’s demo at the website111consistency-tta.github.io/demo. Additionally, a live demo of ConsistencyTTA is available222huggingface.co/spaces/Bai-YT/ConsistencyTTA. Furthermore, the training and inference code is open-sourced333github.com/Bai-YT/ConsistencyTTA.
@@ 2.1 Diffusion Models
Diffusion models [14, 15], known for their diverse and high-quality generations, have rapidly gained popularity across vision and audio generation tasks [10, 16, 3, 17, 18]. In vision, while pixel-level diffusion (e.g., EDM [16]) excels in generating small images, producing larger images requires LDMs [10] as they facilitate the diffusion process within a latent space. In the audio domain, while some works considered autoregressive models [8] or Mel-space diffusion [9], LDMs have emerged as the dominant TTA approach [1, 2, 3, 4, 5, 6, 7].
@@ 4.2 Main Evaluation Results
All diffusion baseline models use 200 inference steps following [3, 1], each step needing two noise estimations due to CFG, summing to 400 network queries per generation. Hence, we conclude that ConsistencyTTA reduces the U-Net queries by a factor of 400 with a minimum performance drop.
@@ 4.4 Audio Generation Diversity
ConsistencyTTA produces diverse generations as do diffusion models. Different random seeds (different initial Gaussian embeddings at $t=T$) produce noticeably different audio. To demonstrate, we present the generated waveforms from the first 50 AudioCaps test prompts with four different seeds at the website444consistency-tta.github.io/diversity. We display the corresponding spectrograms, along with quantitative generation diversity analyses, in Section A.5.
@@ A.2 Real-World Inference Computing Time Comparison
On an Nvidia A100 GPU, generating from all 882 AudioCaps test prompts requires 2.3 minutes with our consistency model. The default TANGO model needs 168 minutes (73 minutes with the smaller 557M U-Net), 72 times as long compared with our consistency model. Note that the 200-step default inference schedule is shared among multiple diffusion-based TTA methods [1, 3], and thus, this TANGO inference time is representative. Moreover, our consistency model can run on a standard laptop computer, only taking 76 seconds to generate 50 ten-second audio clips, averaging 9.1 seconds per minute-generation. I.e., ConsistencyTTA enables on-device audio generation. In contrast, the default TANGO requires 68 seconds per minute-generation on a state-of-the-art A100 GPU.
@@ A.5 More Generation Diversity Evidences
The Mel spectrograms in Figure 3 can also be used to evaluate generation diversity from a quantitative perspective. Specifically, we normalize each spectrogram to have a range of $[0,1]$. Then, for each prompt and each entry of the spectrogram matrix, we calculate the standard deviation across different seeds, resulting in a “standard deviation matrix” with the same shape as the Mel spectrogram. Finally, we average all entries in all “standard deviation matrices”, producing a single number that represents the Mel spectrogram diversity. This number is $0.106$, again demonstrating non-trivial generation diversity.
@@ B.5 Evaluation Details
While the maximal audio length of the AudioCaps dataset is 10.00 seconds and the U-Net module of the TTA models is trained to generate 10.00-second latent audio representations, the HiFi-GAN vocoder produces 10.24-second audio, with the final 0.24 seconds empty. We observe that this mismatch negatively leads to underestimation in generation quality. To this end, when calculating the objective metrics in Tables 1 and 2, we truncate the generated audio to 9.70 seconds (the ground-truth reference waveforms are kept as-is). For $\text{CLAP}_{\text{A}}$ and $\text{CLAP}_{\text{T}}$ calculations, we use the CLAP checkpoint from [40] trained on LAION-Audio-630k [40], AudioSet [41], and music.
The human evaluation results in Table 1 are based on 20 evaluators each rating 25 audio clips per model, forming 500 samples per model. All AudioCaps captions are in English, and all evaluators are proficient in English, using it as their main business language. For each evaluator, the three models and the ground truth use the same set of prompts. Different evaluators are assigned with different prompts and audio clips. Each evaluator rates each audio on a scale of 1 to 5, with rating criteria defined in the evaluation form. To ensure evaluation fairness, the model type generating each waveform is not disclosed to the evaluator, and the generations of the models are shuffled. We find it extremely challenging for a human to distinguish the outputs from the three generative models, with many ground truth waveforms also indistinguishable. An example evaluation form is available at consistency-tta.github.io/evaluation.